Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day08


Crew 100A Log Book for February 19, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: N/A

Brief Narrative of Field Mission Results: Unfortunately it was raining during all day at the MDRS, therefore we did not engage in any EVA. During the morning the time was used to conduct the habitability experiments and follow the MUSK telescope course.

The afternoon was reserved to fixing a problem with the electrical system, a flat tire on the Opportunity ATV and preparation of future experiments. It was time for us to document the work done so far, analyze the first results and prepare a forthcoming EVA with both the drone and the rover to the sites P2,  N38.40746 W110.79280   surface 1382 m and P3,  N38.40737 W110.7961  surface  1375 m. The goal is to take pictures of both these sites and determine the reason why they show such different values of biota and amino acid content.

EVA Data/Interpretations: N/A

Engineering/Hab Maintenance: Flat tire of the Opportunity ATV fixed. electrical problem in the Lab solved.

Plans for Tomorrow: Finish the Tiger mission with an EVA to same location of Feb 18th with astronauts collecting samples for analysis at the laboratory of the Habitat. Continue with experiments: habitability, and EVA to test stress level in the afternoon.

Reports scheduled for today: Engineering, Commander, Journalist

Inventory: None

Support Requested: Refill water tank in trailer

Miscellaneous: N/A


Commander's Report

Ivo Ferreira Reporting

On the 7th day it rained on Mars.

Looks like we are suceeding on terraforming this planet! Today it rained! and not just a bit, it rained the whole day.

Today we woke up without Jeffrey who left the station at around 6am to go back to reality. Yesterday we joked about him coming back to the station after losing his flight and... our predictions were correct... he really missed the flight. Unfortunately he had way too much miles on his account so they booked him for the next flight.

Since it was raining, we decided to wait until the afternoon to finish our Tiger mission (the one we ended with flash lights yesterday).

Certainly the rain would have finished on the afternoon... We used our time to analyze some of the data we already have, conduct as much experiments for Ayako as possible and follow the course to operate the Musk telescope.

We were now in the afternoon and it continued to rain so we decided to start preparing another mission for the SALM (Site Analogue Lune Mars) Group in La Reunion, fix an electrical problem in the station and a flat tire on one of the ATVs.

Water kept falling from sky, mud kept entering at the station, the mood needed to be boosted. Time for our first cake (or that was our plan), an "amazing" chocolate-walnut-whatever-else-we-had cake whose recipe we will never reveal.

The light was gone and finally the rain stopped, what a timing...

Hopefully tomorrow the terraforming process gives us a break and we can finish the Tiger mission.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.78

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.55

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 11 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 18

· Trailer Potable Water Tank Level (inches from bottom): 1/10 Full

· Hab Potable Water Tank Level (inches from bottom): 7

· Potable Water Meter Reading: 423121

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: 1 tire is flat. We will check this tomorrow.

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 73

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Journalist Report

Romain Benchenafi Reporting

Today we have been stuck inside the station because of the rain. So we occupied ourselves with a lot of personal work and some even tried to improve their cooking skills. The cooking mission was a real success so we have had the pleasure to taste an excellent cake!